# Márton Benedek 1122    Péter Biró 2211    Gergely Csáji 22 Matthew Johnson 33

### v2

@@ Abstract
In kidney exchange programmes (KEP) patients may swap their incompatible donors leading to cycles of kidney transplants. Recently, countries try to merge their national patient-donor pools leading to international KEPs (IKEPs). As shown in the literature, long-term stability of an IKEP can be achieved through a credit-based system. In each round, every country is prescribed a “fair” initial allocation of kidney transplants. The initial allocation, which we obtain by using solution concepts from cooperative game theory, is adjusted by incorporating credits from the previous round, yielding the target allocation. The goal is to find, in each round, an optimal solution that closely approximates this target allocation. There is a known polynomial-time algorithm for finding an optimal solution that lexicographically minimizes the country deviations from the target allocation if only $2$-cycles (matchings) are permitted. In practice, kidney swaps along longer cycles may be performed. However, the problem of computing optimal solutions for maximum cycle length $\ell$ is NP-hard for every $\ell\geq 3$. By contrast, the problem is polynomial time solvable once we allow unbounded cycle length. However, in contrast to the case where $\ell=2$, we show that for $\ell=\infty$, lexicographical minimization is only polynomial-time solvable under additional conditions (assuming ${\sf P}\neq{\sf NP}$). Nevertheless, the fact that the optimal solutions themselves can be computed in polynomial time if $\ell=\infty$ still enables us to perform a large scale experimental study for showing how stability and total social welfare are affected when we set $\ell=\infty$ instead of $\ell=2$. omputational complexity, cooperative game theory, partitioned permutation game, international kidney exchange
Keywords:
c
@@ 1 Introduction
We introduce a new class of cooperative games called partitioned permutation games, which are closely related to the known classes of permutation games [40] and partitioned matching games [14]. Partitioned matching games and partitioned permutation games have immediate applications in international kidney exchange. Before defining these games, we first explain this application area.
@@
Kidney Exchange. The most effective treatment for kidney failure is transplanting a kidney from a deceased or living donor, with better long-term outcomes in the latter case. However, a kidney from a family member or friend might be medically incompatible and could easily be rejected by the patient’s body. Therefore, many countries run national Kidney Exchange Programmes (KEPs) [13]. In a KEP, all patient-donor pairs are placed together in one pool. If for two patient-donor pairs $(p,d)$ and $(p^{\prime},d^{\prime})$, it holds that $d$ and $p$ are incompatible, as well as $d^{\prime}$ with $p^{\prime}$, then $d$ could donate a kidney to $p^{\prime}$, and $d^{\prime}$ could donate a kidney to $p$. This is a $2$-way exchange.
@@
We now generalize a $2$-way exchange. We model a pool of patient-donor pairs as a directed graph $G=(V,A)$ (the compatibility graph) in which each vertex of $V$ is a patient-donor pair, and $A$ consists of every arc $(u,v)$ such that the donor of $u$ is compatible with the patient of $v$. In a directed cycle $C=u_{1}u_{2}\ldots u_{k}u_{1}$, for some $k\geq 2$, the kidney of the donor of $u_{i}$ could be given to the patient of $u_{i+1}$ for every $i\in\{1,\ldots,k+1\}$, with $u_{k+1}:=u_{1}$. This is a $k$-way exchange using the exchange cycle $C$. To prevent exchange cycles from breaking (and a patient from losing their willing donor), hospitals perform the $k$ transplants in a $k$-way exchange simultaneously. Hence, KEPs impose a bound $\ell$, called the exchange bound, on the maximum length (number of edges) of an exchange cycle, typically $2\leq\ell\leq 5$. An $\ell$-cycle packing of $G$ is a set ${\mathcal{C}}$ of directed cycles, each of length at most $\ell$, that are pairwise vertex-disjoint; if $\ell=\infty$, we also say that ${\mathcal{C}}$ is a cycle packing. The size of ${\mathcal{C}}$ is the number of arcs that belong to a cycle in ${\mathcal{C}}$, or equivalently, the sum of the lengths of the cycles of ${\cal C}$.
@@
KEPs operate in rounds. A solution for round $r$ is an $\ell$-cycle packing in the corresponding compatibility graph $G^{r}$. The goal is to help as many patients as possible in each round. Hence, to maximize the number of transplants in round $r$, we seek an optimal solution, that is, a maximum $\ell$-cycle packing of $G^{r}$, i.e., one that has maximum size. After round $r$, some patients will have received a kidney or died, while other patient-donor pairs may have arrived. This results in a new compatibility graph $G^{r+1}$ for round $r+1$.
@@
The main computational issue for KEPs is how to find an optimal solution in each round. If $\ell=2$, we can transform, by keeping the “double” arcs, a compatibility graph $G$ into an undirected graph $D=(V,E)$ as follows. For every $u,v\in V$, we have $uv\in E$ if and only if $(u,v)\in A$ and $(v,u)\in A$. It then remains to compute a maximum matching in $D$, which can be done in polynomial time. If we set $\ell=\infty$, a well-known trick works (see e.g. [1]). We transform $G$ into a bipartite graph $H$ with partition classes $V$ and $V^{\prime}$, where $V^{\prime}$ is a copy of $V$. For each $u\in V$ and its copy $u^{\prime}\in V^{\prime}$, we add the edge $uu^{\prime}$ with weight 00. For each $(u,v)\in A$, we add the edge $uv$ with weight $1$. Now it remains to find in polynomial time a maximum weight perfect matching in $H$. However, for any constant $\ell\geq 3$, the complexity changes, as shown by Abraham, Blum and Sandholm [1].
Theorem 1.1 ([1])
@@
International Kidney Exchange. As merging pools of national KEPs leads to better outcomes, many countries today work together with the aim of forming an international KEP (IKEP), e.g. Austria and the Czech Republic [18]; Denmark, Norway and Sweden; and Italy, Portugal and Spain [43]. Apart from ethical, legal and logistical issues (all beyond our scope), there is now a new and highly non-trivial issue that needs to be addressed: How can we ensure long-term stability of an IKEP? If countries are not treated fairly, they may leave the IKEP. In a worst-case scenario it could even happen that in the end each country returns to their own national KEP again.
Example. Let $G$ be the compatibility graph from Figure 1. Then a total of five kidney transplants is possible (only) if we choose the solution of size $5$ that consists of the single exchange cycle $C=abdeca$. We have that patient-donor pairs $a,b,c$ belong to country $1$, whereas patient-donor pairs $d$ and $e$ belong to country $2$, and patient-donor pair $f$ belongs to country $3$. So using this solution, countries $1$, $2$ and $3$ receive three, two and zero kidney transplants, respectively. Note that without cooperation only country $2$ can receive kidney transplants in this example.∎
@@
Figure 1: A permutation game $(N,v)$ of width $c=3$ defined on a graph $G=(V,A)$. Note that $N=\{1,2,3\}$ and $V=V_{1}\cup V_{2}\cup V_{3}$ with $V_{1}=\{a,b,c\}$, $V_{2}=\{d,e\}$ and $V_{3}=\{f\}$, as indicated by the dotted circles. Also note that $v(N)=v(\{1,2\})=5$, as indicated by the thick edges, while $v(\{2\})=2$ and $v(\{1\})=v(\{3\})=0$.
@@
Cooperative Game Theory considers fair distributions of joint profit if all parties involved collaborate (the notion of fairness depends on context). Before describing its role in our setting, we first give some terminology. A (cooperative) game is a pair $(N,v)$, where $N$ is a set of $n$ players and $v:2^{N}\to\mathbb{R}$ is a value function with $v(\emptyset)=0$. A subset $S\subseteq N$ is a coalition. If for every possible partition $(S_{1},\ldots,S_{r})$ of $N$ it holds that $v(N)\geq v(S_{1})+\cdots+v(S_{r})$, then players will benefit most by forming the grand coalition $N$. The problem is then how to fairly distribute $v(N)$ amongst the players of $N$. An allocation is a vector $x\in\mathbb{R}^{N}$ with $x(N)=v(N)$ (we write $x(S)=\sum_{p\in S}x_{p}$ for $S\subseteq N$). A solution concept prescribes a set of fair allocations for a game $(N,v)$.
@@
The solution concepts considered in this paper are the Shapley value, Banzhaf value (two variants) nucleolus, tau value, benefit value, contribution value (all of which are defined in Section 4) and the core. They all prescribe a unique allocation except for the core, which consists of all allocations $x\in\mathbb{R}^{N}$ with $x(S)\geq v(S)$ for every $S\subseteq N$. Core allocations ensure $N$ is stable, as no subset $S$ will benefit from forming their own coalition. But a core may be empty.
@@
We now define some relevant games. For a directed graph $G=(V,A)$ and subset $S\subseteq V$, we let $G[S]=(S,\{(u,v)\in A\;|\;u,v\in S\})$ be the subgraph of $G$ induced by $S$. An $\ell$-permutation game on a directed graph $G=(V,A)$ is the game $(N,v)$, where $N=V$ and for $S\subseteq N$, the value $v(S)$ is the maximum size of an $\ell$-cycle packing of $G[S]$. Two special cases are well studied. We obtain a matching game if $\ell=2$, which may have an empty core (e.g. when $G$ is a triangle on vertices $a,b,c$ with arcs $(a,b)$, $(b,a)$, $(a,c)$, $(c,a)$, $(b,c)$, $(c,b)$) and a permutation game if $\ell=\infty$, whose core is always nonempty [40].
@@
To ensure IKEP stability, we use the model of Klimentova et al. [32], which is a credit-based system. For round $r\geq 1$, let $G^{r}$ be the compatibility graph with associated game $(N^{r},v^{r})$; let $y^{r}$ be the initial allocation (as prescribed by some solution concept ${\mathcal{S}}$); and let $c^{r}:N^{r}\to\mathbb{R}$ be a credit function, which satisfies $\sum_{p\in N{{}^{r}}}c^{r}_{p}=0$; if $r=1$, we set $c^{r}\equiv 0$. For $p\in N$, we set $x^{r}_{p}:=y^{r}_{p}+c^{r}_{p}$ to obtain the target allocation $x^{r}$ for round $r$; note that $x^{r}$ is indeed an allocation, as $y^{r}$ is an allocation and $\sum_{p\in N}c_{p}^{r}=0$). We choose some maximum $\ell$-cycle packing ${\mathcal{C}}$ of $G^{r}$ as optimal solution for round $r$ (out of possibly exponentially many optimal solutions). Let $s_{p}({\mathcal{C}})$ be the number of kidney transplants for patients in country $p$ (with donors both from $p$ and other countries). For $p\in N$, we set $c^{r+1}_{p}:=x^{r}_{p}-s_{p}({\mathcal{C}})$ to get the credit function $c^{r+1}$ for round $r+1$ (note that $\sum_{p\in N}c_{p}^{r+1}=0$). For round $r+1$, a new initial allocation $y^{r+1}$ is prescribed by ${\mathcal{S}}$ for the associated game $(N^{r+1},v^{r+1})$. For every $p\in N$, we set $x_{p}^{r+1}:=y_{p}^{r+1}+c_{p}^{r+1}$, and we repeat the process.
Apart from specifying the solution concept ${\mathcal{S}}$, we must also determine how to choose in each round a maximum $\ell$-cycle packing ${\mathcal{C}}$ (optimal solution) of the corresponding compatibility graph $G$. We will choose ${\mathcal{C}}$, such that the vector $s({\mathcal{C}})$, with entries $s_{p}({\mathcal{C}})$, is closest to the target allocation $x$ for the round under consideration.
To explain our distance measures, let $|x_{p}-s_{p}({\mathcal{C}})|$ be the deviation of country $p$ from its target $x_{p}$ if ${\mathcal{C}}$ is chosen out of all optimal solutions. We order the deviations $|x_{p}-s_{p}({\mathcal{C}})|$ non-increasingly as a vector $d({\mathcal{C}})=(|x_{p_{1}}-s_{p_{1}}({\mathcal{C}})|,\dots,|x_{p_{n}}-s_{p_{n}}({\mathcal{C}})|)$. We say ${\mathcal{C}}$ is strongly close to $x$ if $d({\mathcal{C}})$ is lexicographically minimal over all optimal solutions. If we only minimize $d_{1}({\mathcal{C}})=\max_{p\in N}\{|x_{p}-s_{p}({\mathcal{C}})|\}$ over all optimal solutions, we obtain a weakly close optimal solution. If an optimal solution is strongly close, it is weakly close, but the reverse might not be true. Both measures have been used, see below.
@@
Benedek et al. [9] proved the following theorem for $\ell=2$ (the “matching” case); in contrast, Biró et al. [14] showed that it is NP-hard to find a weakly close maximum matching even for $|N|=2$ once the games are defined on edge-weighted graphs.
Theorem 1.2 ([9])
@@
Benedek et al. [9] used the algorithm of Theorem 1.2 to perform simulations for up to 15 countries for $\ell=2$. As initial allocations, they used the Shapley value, nucleolus, benefit value and contribution value, with the Shapley value yielding the best results. Afterwards, Benedek et al. [10] extended the experimental results from [9] for $\ell=2$ by also considering two variants of the Banzhaf value and the tau value. It turned out that the Banzhaf value variants behaved slightly better than the Shapley value. The good performance of the Shapley value in [9, 10] is in line with the simulation results of Klimentova et al. [32] and Biró et al. [12] for $\ell=3$. Due to Theorem 1.1, the simulations in [12, 32] are for up to four countries and use weakly close optimal solutions.
For $\ell=2$, we refer to [39] for an alternative model based on so-called selection ratios using lower and upper target numbers. IKEPs have also been modelled as non-cooperative games in the consecutive matching setting. In this setting, each round consists of two phases: national pools in phase 1 and a merged pool for unmatched patient-donor pairs in phase 2; see [19, 20, 38] for some results in this setting.
Fairness (versus optimality) issues are also studied for national KEPs, in particular in the US. The US is different from Europe, the setting we consider, as hospitals in the US are more independent and are given positive credits to register their easy-to-match patient-donor pairs to one of the three national KEPs, and negative credits for registering their hard-to-match pairs. The US setting is extensively studied [2, 3, 4, 17, 28, 42], but beyond the scope of our paper.

### v3

# Márton Benedek    Péter Biró    Gergely Csáji Matthew Johnson Daniël Paulusma    Xin Ye

@@ Abstract
In kidney exchange programmes, patients with incompatible donors obtain kidneys via cycles of transplants. Countries may merge their national patient-donor pools to form international programmes. To ensure fairness, a credit-based system is used: a cooperative game-theoretic solution concept prescribes a “fair” initial allocation, which is adjusted with accumulated credits to form a target allocation. The objective is to maximize the number of transplants while staying close to the target allocation. When only $2$-cycles are permitted, a solution that lexicographically minimizes deviations from the target can be found in polynomial time. However, even the problem of maximizing the number of transplants is NP-hard for larger upper bounds on cycle length. This latter problem is tractable when cycle lengths are not bounded. We formalize this setting via a new class of cooperative games called partitioned permutation games, and prove that computing an optimal solution that is lexicographically closest to the target allocation is NP-hard. We give a randomized XP-time algorithm for solve this problem exactly. We present an experimental study, simulating programmes with up to 10 countries. Allowing unbounded cycle lengths increases the number of transplants by up to 46% compared to $2$-cycles. Using credits and selecting lexicographically closest solutions yields low total relative deviation (below 2% for all fairness notions). Among the seven fairness notions tested, a modified Banzhaf value performs best in balancing fairness and efficiency, achieving average deviations below 0.65%. Lexicographic minimization from the target allocation leads to significantly (3636-5656%) smaller average deviations than minimizing maximum difference only.
omputational complexity, cooperative game theory, partitioned permutation game, international kidney exchange
@@ 1 Introduction
We introduce a new class of cooperative games called partitioned permutation games, which are closely related to the known classes of permutation games [TPPR84] and partitioned matching games [BBKPP]. Partitioned matching games and partitioned permutation games have immediate applications in international kidney exchange. Before defining these games, we first explain this application area.
@@
Kidney Exchange. The most effective treatment for kidney failure is transplanting a kidney from a deceased or living donor, with better long-term outcomes in the latter case. However, a kidney from a family member or friend might be medically incompatible and could easily be rejected by the patient’s body. Therefore, many countries run national Kidney Exchange Programmes (KEPs); see [Bi_etal2019, Bi_etal2021] and the recent survey [BCDMP25]. In a KEP, all patient-donor pairs are placed together in one pool. If for two patient-donor pairs $(p,d)$ and $(p^{\prime},d^{\prime})$, it holds that $d$ and $p$ are incompatible as well as $d^{\prime}$ with $p^{\prime}$, but $d$ and $p^{\prime}$ are compatible as well as $d^{\prime}$ with $p$, then $d$ could donate a kidney to $p^{\prime}$, and $d^{\prime}$ could donate a kidney to $p$. This is a $2$-way exchange.
@@
We can generalize a $2$-way exchange as follows. We model a pool of patient-donor pairs as a directed graph $G=(V,A)$ (the compatibility graph) in which each vertex of $V$ is a patient-donor pair, and $A$ consists of every arc $(u,v)$ such that the donor of $u$ is compatible with the patient of $v$. In a directed cycle $C=u_{1}u_{2}\ldots u_{k}u_{1}$, for some $k\geq 2$, the kidney of the donor of $u_{i}$ could be given to the patient of $u_{i+1}$ for every $i\in\{1,\ldots,{k}\}$, with $u_{k+1}:=u_{1}$. This is a $k$-way exchange using the exchange cycle $C$.
To prevent exchange cycles from breaking (and a patient from losing their willing donor), hospitals perform the $k$ transplants in a $k$-way exchange simultaneously. Hence, KEPs impose a bound $\ell$, called the exchange bound, on the maximum length (number of arcs) of an exchange cycle, typically $2\leq\ell\leq 5$. An $\ell$-cycle packing of $G$ is a set $\mathcal{C}$} of directed cycles, each of length at most $\ell$, that are pairwise vertex-disjoint; if $\ell=\infty$, we also say that $\mathcal{C}$} is a cycle packing. The size of $\mathcal{C}$} is is the sum of the lengths of the cycles of $\cal C$}.
@@
KEPs operate in rounds. A solution for round $r$ is an $\ell$-cycle packing in the corresponding compatibility graph $G^{r}$. The goal is to help as many patients as possible in each round. Hence, to maximize the number of transplants in round $r$, we seek an optimal solution, that is, a maximum $\ell$-cycle packing of $G^{r}$, i.e., one that has maximum size. After round $r$, some patients will have received a kidney or died (or have left the pool for some other reason), while other patient-donor pairs may have arrived. This results in a new compatibility graph $G^{r+1}$ for round $r+1$. We have now finished the description of the basic KEP model, as used in this paper. As discussed by Biró et al. [Bi_etal2019], there exist many variants and extensions, some of which we will discuss in more detail later.
@@
An important computational issue for KEPs is how to find an optimal solution in each round. If $\ell=2$, we can apply a well-known transformation (see e.g. [RSU05]) that only keeps the “double” arcs of a compatibility graph $G$. Namely we transform $G$ into an undirected graph ${\hat{G}}=(V,E)$ as follows. For every $u,v\in V$, we have $uv\in E$ if and only if $(u,v)\in A$ and $(v,u)\in A$. It then remains to compute a maximum matching in $\hat{G}$, which can be done in polynomial time [Ed65a]. If we set $\ell=\infty$, another well-known trick works (see e.g. [ABS07]). We transform $G$ into a bipartite graph $H$ with partition classes $V$ and $V^{\prime}$, where $V^{\prime}$ is a copy of $V$. For each $u\in V$ and its copy $u^{\prime}\in V^{\prime}$, we add the edge $uu^{\prime}$ with weight 0. For each $(u,v)\in A$, we add the edge $uv$ with weight $1$. Now it remains to find in polynomial time a maximum weight perfect matching in $H$. However, for any constant $\ell\geq 3$, the complexity changes, as shown by Abraham, Blum and Sandholm [ABS07].
Theorem 1.1([ABS07])
@@
International Kidney Exchange. As merging pools of national KEPs leads to better outcomes, many countries today work together with the aim of forming an international KEP (IKEP), e.g. Austria and the Czech Republic [Bo_etal17]; Denmark, Norway and Sweden; and Italy, Portugal and Spain [Va_etal19]. Apart from ethical, legal and logistical issues (all beyond our scope), there is now a new and highly non-trivial issue that needs to be addressed: How can we ensure long-term stability of an IKEP? If countries are not treated fairly, they may leave the IKEP. In a worst-case scenario it could even happen that in the end each country returns to their own national KEP again.
Example 1. Let $G$ be the compatibility graph from Figure 1. Then a total of five kidney transplants is possible (only) if we choose the solution of size $5$ that consists of the single exchange cycle $C=abdeca$. We have that patient-donor pairs $a,b,c$ belong to country $1$, whereas patient-donor pairs $d$ and $e$ belong to country $2$, and patient-donor pair $f$ belongs to country $3$. So using this solution, countries $1$, 22 and $3$ receive three, two and zero kidney transplants, respectively. Note that without cooperation only country $2$ can receive kidney transplants in this example.∎
@@
Figure 1: A partitioned permutation game $(N,v)$ of width $c=3$ defined on a graph $G=(V,A)$. Note that $N=\{1,2,3\}$ and $V=V_{1}\cup V_{2}\cup V_{3}$ with $V_{1}=\{a,b,c\}$, $V_{2}=\{d,e\}$ and $V_{3}=\{f\}$, as indicated by the dotted circles. Recall that $G$ has exactly one maximum cycle packing $\cal C$}, which consists of the single exchange cycle $C=abdeca$, as indicated by the thick edges. Selecting $\cal C$} yields the allocation $x=(3,2,0)$. We note that $v(N)=v(\{1,2\})=5$, while $v(\{2\})=2$ and $v(\{1\})=v(\{3\})=v(\{1,3\})=v(\emptyset)=0$ and $v(\{2,3\})=2$. It can be readily checked that for all $S\subseteq N$, $x(S)\geq v(S)$. Hence, $x$ belongs to the core of $(N,v)$.
Cooperative Game Theory considers fair distributions of joint profit if all parties involved collaborate (the notion of fairness depends on context). Before describing its role in our setting, we first give some basic terminology on cooperative game theory; we refer to [Peters2008] for more details.
@@
A (cooperative) game is a pair $(N,v)$, where $N$ is a set of $n$ players and $v:2^{N}\to\mathbb{R}$ is a value function with $v(\emptyset)=0$. A subset $S\subseteq N$ is a coalition. If for every possible partition $(S_{1},\ldots,S_{r})$ of $N$ it holds that $v(N)\geq v(S_{1})+\cdots+v(S_{r})$, then players will benefit most by forming the grand coalition $N$. The problem is then how to fairly distribute $v(N)$ amongst the players of $N$. An allocation is a vector $x\in\mathbb{R}^{N}$ with $x(N)=v(N)$ (we write $x(S)=\sum_{p\in S}x_{p}$ for $S\subseteq N$). In this context, the notion of fairness is defined by a solution concept, which prescribes a set of fair allocations for a game $(N,v)$. Each solution concept has its own fairness properties; we discuss these properties in detail in Section 4.2.
@@
The solution concepts considered in this paper are the Shapley value, Banzhaf value (two variants) nucleolus, tau value, benefit value, contribution value (all of which are defined in Section 4.2) and the core. They all prescribe a unique allocation except for the core, which consists of all allocations $x\in\mathbb{R}^{N}$ with $x(S)\geq v(S)$ for every $S\subseteq N$. Core allocations ensure $N$ is stable in the sense that no subset $S$ will benefit from forming their own coalition. However, the core of a cooperative game may be empty.
@@
We now define some relevant games. For a directed graph $G=(V,A)$ and subset $S\subseteq V$, we let $G[S]=(S,\{(u,v)\in A\;|\;u,v\in S\})$ be the subgraph of $G$ induced by $S$. An $\ell$-permutation game on a directed graph $G=(V,A)$ is the game $(N,v)$, where $N=V$ and for $S\subseteq N$, the value $v(S)$ is the maximum size of an $\ell$-cycle packing of $G[S]$. Two special cases are well studied. We obtain a matching game if $\ell=2$, which may have an empty core (e.g. when $G$ is a triangle on vertices $a,b,c$ with arcs $(a,b)$, $(b,a)$, $(a,c)$, $(c,a)$, $(b,c)$, $(c,b)$, see also Figure 2) and a permutation game if $\ell=\infty$, whose core is always nonempty [TPPR84].
In the remainder, we differentiate between the sets $N$ (of players in the game) and $V$ (of vertices in the underlying graph $G$). That is, we associate each player $i\in N$ with a distinct subset $V_{i}$ of $V$. A partitioned $\ell$-permutation game on a directed graph $G=(V,A)$ with a partition $(V_{1},\ldots,V_{n})$ of $V$ is the game $(N,v)$, where $N=\{1,\ldots,n\}$, and for $S\subseteq N$, the value $v(S)$ is the maximum size of an $\ell$-cycle packing of $G[\bigcup_{i\in S}V_{i}]$. We obtain a partitioned matching game [BBJPX23, BBKPP] if $\ell=2$, and a partitioned permutation game if $\ell=\infty$. The width of $(N,v)$ is the width of $(V_{1},\ldots,V_{n})$, which is defined as $c=\max\{|V_{i}|\;|\;1\leq i\leq n\}$. We refer again to Figure 1 for an example.
To ensure IKEP stability, we use the model of Klimentova et al. [KNPV20], which is a credit-based system. For round $r\geq 1$, let $G^{r}$ be the compatibility graph with associated game $(N,v^{r})$; let $y^{r}$ be the initial allocation (as prescribed by some solution concept $\mathcal{S}$}); and let $c^{r}:N\to\mathbb{R}$ be a credit function, which satisfies $\sum_{p\in N}c^{r}_{p}=0$; if $r=1$, we set $c^{r}\equiv 0$. For $p\in N$, we set $x^{r}_{p}:=y^{r}_{p}+c^{r}_{p}$ to obtain the target allocation $x^{r}$ for round $r$; note that $x^{r}$ is indeed an allocation, as $y^{r}$ is an allocation and $\sum_{p\in N}c_{p}^{r}=0$). We choose some maximum $\ell$-cycle packing $\mathcal{C}$} of $G^{r}$ as optimal solution for round $r$ (out of possibly exponentially many optimal solutions). For round $r$, let $s_{p}({\mathcal{C}})^{{r}}$ be the number of kidney transplants for patients in country $p$ (with donors both from $p$ and other countries). For $p\in N$, we set $c^{r+1}_{p}:=x^{r}_{p}-s_{p}({\mathcal{C}}^{{r}})$ to get the credit function $c^{r+1}$ for round $r+1$ (note that $\sum_{p\in N}c_{p}^{r+1}=0$). For round $r+1$, a new initial allocation $y^{r+1}$ is prescribed by $\mathcal{S}$} for the associated game $(N,v^{r+1})$. For every $p\in N$, we set $x_{p}^{r+1}:=y_{p}^{r+1}+c_{p}^{r+1}$, and we repeat the process.
Apart from specifying the solution concept $\mathcal{S}$}, we must also determine how to choose in each round a maximum $\ell$-cycle packing $\mathcal{C}$} (optimal solution) of the corresponding compatibility graph $G$. We will choose $\mathcal{C}$}, such that the vector $s({\mathcal{C}})$, with number of kidney transplants entries $s_{p}({\mathcal{C}})$, is closest to the target allocation $x$ for the round under consideration. To ensure (long-term) stability of an IKEP, we aim to make an IKEP balanced, that is, in each round the goal is to find optimal cycle packings that are close to the target allocations, yielding consistently low (ideally zero) credit values.
To explain our distance measures, let $|x_{p}-s_{p}({\mathcal{C}})|$ be the deviation of country $p$ from its target $x_{p}$ if $\mathcal{C}$} is chosen out of all optimal solutions. We order the deviations $|x_{p}-s_{p}({\mathcal{C}})|$ non-increasingly as a vector
$d({\mathcal{C}})=(|x_{p_{1}}-s_{p_{1}}({\mathcal{C}})|,\dots,|x_{p_{n}}-s_{p_{n}}({\mathcal{C}})|).$
We say that $\mathcal{C}$} is strongly close to $x$ if $d({\mathcal{C}})$ is lexicographically minimal over all optimal solutions. If we only minimize $d_{1}({\mathcal{C}})=\max_{p\in N}\{|x_{p}-s_{p}({\mathcal{C}})|\}$ over all optimal solutions, we obtain a weakly close optimal solution. If an optimal solution is strongly close, then it is weakly close, but the reverse might not be true. Both measures have been used in the literature, as we discuss below after first giving an example.
Example 2. Let $\ell=2$ and consider the partitioned matching game $(N,v)$ from Figure 2, which has width $c=1$. Note that $N=\{1,2,3\}$ and $V=V_{1}\cup V_{2}\cup V_{3}$ with $V_{1}=\{a\}$, $V_{2}=\{b\}$ and $V_{3}=\{c\}$. Moreover, $v(N)=v(\{1,2,3\}=v(\{1,2\})=v(\{1,3\})=v(\{2,3\})=2$, while $v(\{1\})=v(\{2\})=v(\{3\})=v(\emptyset)=0$. Say we are in round 1 and use the Shapley value (see Section 4), which yields the initial allocation $x^{1}=\{\frac{2}{3},\frac{2}{3},\frac{2}{3}\}$. As $c^{1}=(0,0,0)$, we have $y^{1}=x^{1}$ as target allocation. Note that $y^{1}$ is not in the core, because the core of $(N,v)$ is empty, as we observed before. There are three optimal solutions, which each correspond to a cycle packing consisting of exactly one $2$-vertex cycle. By symmetry, each of them is strongly close to $x$. Suppose we select ${\cal C}=\{\{aba\}\}$. Note that $s_{1}({\cal C})=s_{2}({\cal C})=1$ and $s_{3}({\cal C})=0$. Hence, $d({\cal C})=(\frac{1}{3},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{3},\frac{2}{3}})$, and the credit vector $c^{2}$ for round $2$ is $(-\frac{1}{3},-\frac{1}{3},\frac{2}{3})$. ∎
@@
Figure 2: A partitioned matching game $(N,v)$ of width $c=1$ on the graph $G=(V,A)$ with $N=\{1,2,3\}$ and $V=V_{1}\cup V_{2}\cup V_{3}$ for $V_{1}=\{a\}$, $V_{2}=\{b\}$ and $V_{3}=\{c\}$. Note that $v(\{1\})=v(\{2\})=v(\{3\})=v(\emptyset)=0$, whereas $v(N)=v(\{1,2,3\}=v(\{1,2\})=v(\{1,3\})=v(\{2,3\})=2$, implying the core of $(N,v)$ is empty.
@@
Benedek et al. [BBKPP] proved the following theorem for $\ell=2$ (the “matching” case).
Theorem 1.2([BBKPP])
@@
In contrast to Theorem 1.2, Benedek et al. [BBKPP] also showed that the problem becomes NP-hard for edge-weighted graphs, in which the edges $e\in A$ have an associated weight $w(e)$ reflecting the expected utility of the corresponding kidney transplant. Namely, if maximum weight matchings are taken as optimal solutions for edge-weighted graphs, then it is NP-hard to find a weakly close optimal solution even for $n=2$.
Benedek et al. [BBKP22] used the algorithm of Theorem 1.2 to perform simulations for up to 15 countries for $\ell=2$. As initial allocations, they used the Shapley value, nucleolus, benefit value and contribution value, with the Shapley value yielding the best results. Afterwards, Benedek et al. [BBPY24] extended the experimental results from [BBKP22] for $\ell=2$ by also considering two variants of the Banzhaf value and the tau value. It turned out that the Banzhaf value variants behaved slightly better than the Shapley value.
The good performance of the Shapley value in [BBKP22, BBPY24] is in line with the simulation results of Klimentova et al. [KNPV20] and Biró et al. [BGKPPV20] for $\ell=3$. Due to Theorem 1.1, the simulations in [BGKPPV20, KNPV20] are for up to four countries and use weakly close optimal solutions. We note that the simulations in [BGKPPV20, KNPV20] model IKEPs that allow for non-directed donors. These are altruistic donors that are added to the pool without a corresponding patient. By allowing non-directed donors, optimal solutions for compatibility graphs may include exchange chains, which are directed paths that start from a non-directed donor. This may result in more kidney transplants.
For $\ell=2$, we refer to [STW21] for an alternative model based on so-called selection ratios using lower and upper target numbers. IKEPs have also been modelled as non-cooperative games in the consecutive matching setting. In this setting, each round consists of two phases: national pools in phase 1 and a merged pool for unmatched patient-donor pairs in phase 2; see [CL23, CLPV17, SBS] for some results in this setting.
Finally, fairness (versus optimality) issues are also studied for national KEPs, in particular in the US. The US setting is different from the European setting, which is the setting we consider, for the following reasons. As we explained above, in Europe the treatment of patient-donor pairs is regulated centrally implying that all patient-donor pairs are registered at an (I)KEP, leading to large and diverse patient-donor pair pools with many exchange options and matching runs that are conducted typically every three months. In contrast, in the US the transplant centers (hospitals) are independent. They mostly treat incoming patient-donor pairs in-house, and need to be convinced to register patient-donor pairs to one of the three nationwide KEPs (UNOS, APD, and NKR) which compete against each other (see [AAAFK18]). Therefore, in the US, kidney transplants usually take place immediately after the registration of a new pair. How to incentivize KEPs in the US setting to register all their patient-donors pairs (instead of only the hard-to-match pairs) to a joint pool is a challenging problem that has been widely studied; see e.g. [AFKP15, AR12, AR14, BCHPPV17, HDHSS15, TP15] for a number of different approaches, which were compared to each other in the recent survey [BCDMP25]. In particular, Hajaj at al. [HDHSS15] suggested a credit-based approach, namely to give the nationwide KEPs positive credits for registering easy-to-match pairs and negative credits for registering hard-to-match pairs. Their approach was further analysed by Agarwal et al. [AAAFK18] and has been adapted by the NKR but is very different in nature from our credit-based approach.